OpenStudy (anonymous):

ok i am stuck on this steps 1/6[(ln4)-(ln3)] please help me solve thanks

OpenStudy (lalaly):

thats the question?

OpenStudy (lalaly):

thats right, do u know the rules of logarithms?

OpenStudy (anonymous):

nope i was hoping you could help me out with where i got up to

OpenStudy (lalaly):

ln is the natural log right? so when u have lnx-lny u can write it as ln(x/y)

OpenStudy (lalaly):

so this is 1/6 ln(4/3)

OpenStudy (lalaly):

1) log(mn) = log(m) + log(n) 2) log(m/n) = log(m) – log(n) 3) log(m^n) = n · log(m)

OpenStudy (anonymous):

ok so when i multiply how do i do that?

OpenStudy (lalaly):

im sorry but multiply what?

OpenStudy (anonymous):

1/6 ln (4/3) or is thats it?

OpenStudy (lalaly):

if u mean the answer is ( 1/6) * ln(4/3) on the calculator or u just leave it like that coz ln(4/3) is anumber u cn write it as ln(4/3) / 6